Beth Zion Hospital has received initial certification from the state of Califonia to become a center for liver transplants. The hospital, however, must complete its first 18 transplants under great scrutiny and at no cost to the patients. The very first transplant. Just completed required 28 hours. On the basis of research at the hospital Beth Zion estimates hat it wil have an 80% learning curve. Using able 3 estimate e time it will take to complete: a) the 5th transplant (round your response to two decimal places). hours b) the first 5 transplants fround your response to two decimal places). hours c) the 18th transplant round your response to two decimal places). hours d) the first 18 transplants (round your response to two decimal placos). hours 1: Learning Curve Coefficients 75% 85% Total Time 3.459 8.245 Total Time 3.738 9.716 Unit Number Total Unit Time 513 301 Unit Time 596 Unit 686 508 4.031 11.405 18
Answer #1

By using the learning curve coefficient approach, the following is found: Where N refers to the total labor hours required for the production of Nth unit T1 refers to the number of labor hours required for the production of first unit

C refers to the value as per learning curve coefficents table

A. From learning curve coefficients table, T1 = 28, C = 0.596
TN= 28*0.596 = 16.688 hours
Time required to complete the 5th transplant = 16.688 hours

B. From learning curve coefficients table, T1 = 28, C = 3.738
TN= 28*3.738 = 104.664 hours
Time required to complete the first five transplants = 104.664 hours

C. From learning curve coefficients table, T1 = 28, C = 0.394
TN= 28*0.394 = 104.664 hours
Time required to complete the first five transplants = 11.032 hours

D. From learning curve coefficients table, T1 = 28, C = 9.716
TN= 28*9.716 = 272.048 hours
Time required to complete the first five transplants = 272.048 hours